# Good Noise Makes Good Edits: A Training-Free Diffusion-Based Video Editing with Image and Text Prompts

### v1

@@ Abstract
We propose ImEdit, the first zero-shot, training-free video editing method conditioned on both images and text. The proposed method introduces $\rho$-start sampling and dilated dual masking to construct well-structured noise maps for coherent and accurate edits. We further present zero image guidance, a controllable negative prompt strategy, for visual fidelity. Both quantitative and qualitative evaluations show that our method outperforms state-of-the-art methods across all metrics.
Figure 1: Comparison results of video editing. ImEdit effectively edits source videos using image and text prompts, achieving high fidelity and temporal consistency without any additional training.
@@ 1 Introduction
Diffusion-based video editing methods have shown significant achievements in recent years, enabling precise modifications of objects in videos based on prompts provided by users stablevideo ; videdit ; videoswap ; genvideo ; videop2p ; dreamix ; revideo ; fatezero ; videoanydoor ; tune_a_video ; controlvideo ; make_a_pro ; cutandpaste . Primarily, text-driven video editing models have gained considerable attention for their ability to modify video content while maintaining temporal consistency tokenflow ; videop2p ; fatezero ; tune_a_video . These methods have demonstrated impressive results across a wide range of applications, such as appearance editing stablevideo ; videdit ; magicedit ; videop2p , and style transfer structure ; tokenflow ; rerenderavideo ; controlvideo .
However, text-based editing has inherent limitations, particularly when modifying objects with intricate and unique details, where textual descriptions alone are often insufficient to precisely convey the desired attributes precisely (Fig. 2(a)). Moreover, crafting detailed textual prompts is cumbersome, requiring extensive trial-and-error to achieve the intended modifications. This challenge becomes even more pronounced when the editing method necessitates one-shot finetuning for each edit videop2p ; fatezero ; save . A more efficient and intuitive alternative is image-driven prompting, which provides direct visual guidance for more precise and straightforward modifications.
Despite its advantages, image-driven video editing, particularly training-free approaches, remains largely unexplored. Previous work, Make-A-Protagonist make_a_pro , introduces an approach that trains a text-to-video (T2V) diffusion model conditioned on both text and image inputs via CLIP encoders clip . Similarly, GenVideo genvideo proposes a mask generation approach using an image diffusion model to address shape and size mismatches between the source object in the source video and the target object in the reference image. While effective, these methods require one-shot fine-tuning for each video, making the editing process highly costly and sensitive to variations in input resolution.
To overcome these limitations, a training-free video editing approach that conditions on both image and text prompts is essential. However, a key challenge remains: maintaining temporal consistency while preserving motion dynamics is fundamental to video editing, but becomes particularly difficult in a training-free setting, leading to color shifts, shape distortions, or loss of fine details. As shown in Fig. 2(b), simply integrating an image prompt into an existing T2V model without fine-tuning results in significant quality degradation, particularly in object shape fidelity.
@@
Figure 2: Importance of appropriate integration of image prompts in video editing. Given (a) a source video, we compare (b) a state-of-the-art text-driven video editing method, (c) a naive combination of a pretrained T2V model with an image-guided module, and (d) our proposed method. The text-only approach in (b) struggles to accurately transfer the fine-grained details of the car in the reference image. Simply attaching the image-guided module to a pretrained T2V model, as in (c), produces unnatural results. In contrast, (d) our proposed training-free method precisely transfers the car in the reference image into the video with high fidelity.
Therefore, we propose that well-structured noise maps, generated through a controlled combination of the source video’s latent representation and Gaussian noise, effectively preserve motion consistency even without additional training. These noise maps act as a key component in achieving high-quality video edits. To further refine the noise maps to ensure seamless object transitions, we introduce a dilated dual masking strategy that blends the target object with the original video, which becomes particularly effective in handling shape discrepancies between the source and target objects. We also observe that a simple yet effective enhancement to the existing negative image prompt formulation used in IP-Adapter ipadapter effectively leads to realistic edits without requiring additional training.
Building on these findings, we present ImEdit, the first zero-shot, training-free video editing approach conditioned on both image and text prompts, enabling high-fidelity object replacement while preserving temporal consistency. Our method leverages well-structured craft noise maps for motion preservation, dilated dual masking for seamless object replacement, and zero image guidance for color fidelity. Additionally, Our approach is highly flexible to varying image resolutions, by leveraging a T2V backbone pretrained on diverse resolutions without additional fine-tuning, ensuring robust generalization across different input scales.
As illustrated in Fig. 2(c), our proposed method effectively modifies video objects based on reference images and simple text prompts, eliminating the need for complex prompt engineering. We also evaluate the time efficiency of the proposed method against baseline methods in Fig. 3. Across all tested input video sizes, our method consistently achieves the fastest runtime, demonstrating its advantage as a training-free solution for video editing. Notably, Fig. 1 highlights that our method effectively addresses challenging scenarios, including small source objects (top), significant color differences between the source and target objects (bottom left), and dynamic motion in the source video (bottom right), while maintaining motion consistency throughout the edited results. In contrast, baseline methods often fail to accurately reflect the reference appearance under these settings.
Figure 3: Comparison of runtime across different methods. Results are averaged over three runs for each unique input video size, with standard deviation.
Through extensive quantitative and qualitative comparisons, we demonstrate that our approach outperforms existing state-of-the-art video editing methods in both visual quality and temporal stability. By eliminating the need for fine-tuning, our method is efficient, scalable, and widely applicable across various video editing tasks. In summary, our key contributions are as follows:
  • We propose ImEdit, the first zero-shot, training-free video editing approach that leverages both text and image prompts to seamlessly replace objects while preserving original motion dynamics and temporal consistency.
  • We introduce an effective noise map construction strategy to guide the denoising U-Net for coherent and accurate video edits. As key components, $\rho$-start sampling preserves motion patterns and spatial layout of the source video, while dilated dual masking ensures smooth transitions in appearance and shape.
  • To further enhance color saturation and visual fidelity, we present zero image guidance, a controllable negative image prompt strategy that blends a zero image embedding into the existing negative image guidance.
  • Comprehensive quantitative and qualitative evaluations demonstrate that our method outperforms existing state-of-the-art approaches.
@@
Figure 4: Overview of our proposed training-free approach. ImEdit constructs two levels of noise maps to guide the editing process. First, a rough noise map is constructed from the source video and denoised using a pretrained T2V model with image and text guidance, producing a rough edit. Next, a refined good noise map is constructed and denoised under the same conditions to produce the final video. In both stage, we denoise for $\rho$ steps by employing $\rho$-start sampling. The output more faithfully reflects the input prompts while preserving motion and scene structure.
@@ 2 Methods
We introduce ImEdit, a training-free video editing approach that performs object replacement while preserving the overall coherence of the scene, guided by text or image prompts. Our method is grounded in the careful integration of well-structured noise with appropriately chosen components. Specifically, we consider two key ingredients: the foreground latent, obtained via $\rho$-start sampling, and the background latent, acquired through DDIM inversion of the source video. These latents are blended using a carefully designed mask derived from both the source and target object masks, employing dilated dual masking to ensure smooth transitions and spatial consistency. In addition, we apply zero image guidance, a negative prompt-based strategy, to enhance color saturation and overall visual fidelity in the final output.
@@ 2.1 Preliminaries
Latent diffusion models (LDMs)
In diffusion-based image generation models, LDMs have become a standard approach due to their computational efficiency and ability to generate high-resolution, photorealistic images dalle2 ; stablediffusion ; imagen . These models leverage a pretrained autoencoder vae , consisting of an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$, to map input images into a lower-dimensional latent space. Given an input image ${\bm{x}}$, the clean latent variable is obtained as ${\bm{z}}_{0}=\mathcal{E}({\bm{x}})$. During the forward process, gaussian noise is gradually added to the latent variable over $T$ steps. In the reverse process, a denoising network $\epsilon(\cdot)$ predicts noise at each step to iteratively refine the latent variable. Once the denoising process is complete, the decoder $\mathcal{D}(\cdot)$ transforms the latent variable back into the image space to generate the final output.
DDIM sampling and inversion
[…]
DDIM inversion predicts the sequence of noisy latents $\{{\bm{z}}_{t}\}_{t=1}^{T}$ from a given clean sample ${\bm{z}}_{0}$, reconstructing the corresponding diffusion trajectory without requiring any model retraining ddim ; inverting . We denote this inversion trajectory as DDIM-inv$({\bm{z}}_{0})$.
@@ 2.2 Overview
Our approach achieves this goal through a two-stage editing pipeline, as illustrated in Fig. 4. In the first stage, we construct a rough noise to enable localized object modification. In the second stage, this rough noise is refined into a more well-structured good noise, leading to higher-fidelity editing.
@@ 2.3 How to construct a good noise for video editing?
We first encode the source video $\mathcal{V}_{s}$ into the latent space using a pretrained VAE encoder $\mathcal{E}$. We then initialize the starting noisy latent using $\rho$-start sampling (see Section 2.3.2), a modification of the standard DDIM sampling schedule, which enables better preservation of the source video’s structure and motion. This noisy latent is passed into a pretrained T2V diffusion model $\epsilon_{\theta}$ where we incorporate both image and text guidance using the IP-Adapter ipadapter . The image and text features from CLIP image encoder $\text{CLIP}_{\text{img}}$ and CLIP text encoders $\text{CLIP}_{\text{txt}}$ clip are injected into the each corresponding cross-attention layers of the T2V model as condition embeddings:
@@
During the denoising process with condition embeddings, we employ a foreground-background separation mechanism to perform object-centric modifications while preserving background fidelity, inspired by latent blending latentblending ; blendeddiffusion ; videoswap ; genvideo . Specifically, we split the noisy latent ${\bm{z}}_{t}$ into foreground and background regions using the object segmentation mask ${\bm{M}}$. For the background region, we retrieve the DDIM inversion noise corresponding to the current time step to preserve the background of the source video $\mathcal{V}_{s}$. In contrast, the foreground region is denoised by iteratively updating the noisy latent from the previous time step using a denoising network $\epsilon_{\theta}$, guided by the condition embeddings. This enables the foreground object to be progressively edited to resemble the target object. Formally, the foreground latent ${\bm{z}}_{t}^{\text{{fg}}}$ and background latent ${\bm{z}}_{t}^{\text{{bg}}}$ at time step $t$ are computed as:
[…]
${\bm{z}}^{\text{{bg}}}_{t\texttt{-}1}$ $={\bm{z}}_{t\texttt{-}1},\ \text{with }{\bm{z}}_{t-1}\in\{{\bm{z}}_{t}\}_{t=1}^{\tau_{\nu}}\text{ from }\text{DDIM-inv}({\bm{z}}_{0}).$ (4)
We then blend these two latents using ${\bm{M}}$ via our dilated dual masking (Section 2.3.1), as follows:
@@
where $\odot$ denotes element-wise multiplication. We use different mask $M$ at each stage. Specifically, in the first stage, the dilated source mask from the source video, ${\bm{M}}^{\prime}_{s}\in\mathbb{R}^{F\times C_{s}\times H_{s}\times W_{s}}$, is used. Through the denoising process in the first stage, we obtain a roughly edited video $\mathcal{V}_{r}$, which reflects the general structure and appearance of the target object. However, this rough edit may still retain residual features of the source object, which should be removed for a clean transformation.
@@
Figure 5: Effective noise construction using our dilated dual masking. We blend the foreground and background latents, obtained through our proposed methods such as $\rho$-start sampling and zero image guidance, using a final mask constructed by taking the union of the dilated source and target masks. This enables clean removal of the source object and seamless replacement with the target.
@@ 2.3.1 Dilated dual masking
The effectiveness of our method stems from the careful coordination of structured noise using the proposed dilated dual mask, which plays a crucial role in defining the editable regions within each frame. This mask is central to enabling selective editing of target objects, as illustrated in Fig. 5.
In the first stage, only the source object mask ${\bm{M}}_{s}$ is available, which roughly outlines the foreground region to be edited. However, this mask lacks contextual adaptation to the background, often resulting in sharp edges and unnatural transitions at the object boundaries. Thus, we apply morphological dilation to smooth boundaries and facilitate seamless object replacement. Specifically, a single-iteration morphological dilation is applied to a binary matrix $M$ with a kernel of size $k\times k$. For each pixel position $(p,q)$, the output value is determined as the maximum value within a neighborhood defined by ${\mathcal{N}}_{k}$, where ${\mathcal{N}}_{k}=\{(i,j)\mid-r\leq i,j\leq r\}$ and $r=\lfloor\frac{k}{2}\rfloor$. The dilation operation is formally expressed as follows:
@@
In the second stage, we similarly obtain the dilated rough mask ${\bm{M}}^{\prime}_{r}$ from the rough edits. However, blending foreground and background latents using either the source mask ${\bm{M}}^{\prime}_{s}$ or the rough mask ${\bm{M}}^{\prime}_{r}$ poses the risk that parts of the source video may remain visible or that the generation area for the target object may not be fully covered. To address this, we leverage masks from both the first and second stages. The final mask ${\bm{M}}_{e}$ is computed as the union of the source mask ${\bm{M}}^{\prime}_{s}$ and the rough mask from the first stage ${\bm{M}}^{\prime}_{r}$, as shown in the following equation:
@@
Finally, we construct the good noise ${\bm{z}}^{\prime}_{t}$ using the final mask ${\bm{M}}_{e}$ to guide the final blending process in Eq. 5. This enables the second stage to generate more realistic, artifact-free results, as ${\bm{M}}_{e}$ precisely defines the edited region. In particular, it ensures clean removal of the original object, accurate reconstruction of the new object’s shape, and consistent preservation of the background.

### v2

@@ Abstract
We propose VINO, the first zero-shot, training-free video editing method conditioned on both image and text. Our approach introduces $\rho$-start sampling and dilated dual masking to construct structured noise maps that enable coherent and accurate edits. To further enhance visual fidelity, we present zero image guidance, a controllable negative prompt strategy. Extensive experiments demonstrate that VINO faithfully incorporates the reference image into video edits, achieving strong performance compared to state-of-the-art baselines, all without any test-time or instance-specific training.
@@ 1 Introduction
Diffusion-based video editing methods have shown notable achievements in recent years, enabling precise modifications of objects in videos based on user prompts [stablevideo, videdit, dreamix, revideo, controlvideo, cutandpaste]. Primarily, text-driven video editing models have gained considerable attention for their ability to modify video content while maintaining temporal consistency [tokenflow, videop2p, fatezero, tune_a_video]. These methods have demonstrated impressive results across a wide range of applications, such as appearance editing [stablevideo, videdit, magicedit, videop2p], and style transfer [structure, tokenflow, rerenderavideo, controlvideo].
@@
Figure 2: Importance of appropriate integration of image prompts in video editing. Given (a) a source video, we compare (b) a state-of-the-art text-driven video editing method, (c) a naive combination of a pretrained T2V model with an image-guided module, and (d) our proposed method. The text-only approach in (b) struggles to accurately transfer the structural appearance of the car in the reference image. Simply attaching the image-guided module to a pretrained T2V model, as in (c), produces unnatural results. In contrast, our proposed training-free method in (d) precisely transfers the car in the reference image into the video with high fidelity.
Figure 3: Runtime comparison across video resolutions. We compare VINO with Make-A-Pro [make_a_pro], an image-guided video editing method. Results are averaged over three runs per setting.
However, text-based video editing has inherent limitations, particularly when modifying objects with intricate and unique details, where textual descriptions alone are often insufficient to precisely convey the desired attributes (Fig. 2(a)). Moreover, crafting detailed textual prompts is cumbersome, requiring extensive trial-and-error to achieve the intended modifications. This challenge becomes even more pronounced when the editing method necessitates one-shot finetuning for each edit [videop2p, fatezero, save].
To mitigate the need for costly text prompt engineering, recent work has explored video editing methods guided by both image and text. However, such approaches [genvideo, make_a_pro] suffer from weakened temporal consistency and require per-video finetuning, making them unsuitable for real-world editing scenarios. Other approaches employ additional guidance such as user-defined semantic keypoints alongside image and text prompts [videoanydoor, videoswap] to achieve more precise control. While effective for temporally coherent local editing guided by reference images, these methods require substantial user interaction and additional training, limiting their practicality.
Therefore, we propose VIdeo editing with NOise control (VINO), the first zero-shot, training-free video editing framework conditioned on both image and text prompts. As illustrated in Fig. LABEL:fig:teaser, our method enables localized video edits guided by reference images, without any additional training or cumbersome user interaction.
We leverage recent advances in the image domain to enable high-fidelity video editing without finetuning. In particular, pretrained image encoders and adapters [pbe, chen2024zero, stableviton, ipadapter] have shown success in injecting reference features into cross-attention layers. However, directly connecting these image-based techniques to existing text-to-video (T2V) models produces unnatural results, pasting the reference image onto the video without considering the source video’s layout (Fig. 2(c)). We address this semantic unnaturalness by introducing $\rho$-start sampling inspired by [sdedit], which starts DDPM sampling from intermediate time steps to generate realistic images from stroke paintings. In our strategy, both DDIM sampling and inversion begin from intermediate time steps, preserving the layout and motion of the source video.
While this strategy effectively balances the appearance of the reference and the layout of the source, it still struggles when replacing source objects with targets of significantly different sizes or shapes (Fig. 7(a)). To overcome object size and shape discrepancies, we design VINO with a two-stage dilated dual masking strategy. In the first stage, we apply dilated masking to blend source video latents with editing latents during the sampling process, producing a coarse edited video. In the second stage, this coarse edited video is refined using dilated dual masking, enabling more flexible spatial blending for accurate appearance transfer.
@@
Figure 4: Overview of VINO. Our proposed training-free approach, VINO, adopts a coarse-to-fine scheme to guide the two-stage editing process via strategically designed noise maps. Both stages leverage $\rho$-start sampling with text and image prompts. In the first stage, a rough noise map is constructed from the source video via dilated masking (a-1), which is then denoised by the pretrained T2V model to yield a rough edit (b-1). In the second stage, a refined good noise map is derived from the rough edit via dilated dual masking (a-2). Finally, the model produces the final output based on this good noise map (b-2), faithfully reflecting the prompts while preserving motion and structure.
We further present zero-image guidance to address occasional color desaturation. This controllable negative prompt strategy steers generation away from the zero-image direction, thereby improving color fidelity. Consequently, our method achieves high visual quality and temporal consistency (Fig. 2(d)), while offering clear efficiency advantages over methods requiring finetuning (Fig. 3).
In summary, our key contributions are:
  • We propose VINO, the first zero-shot video editing method using both text and image prompts, preserving motion and temporal consistency without training.
  • We introduce a noise map construction strategy including $\rho$-start sampling to preserve the motion patterns and spatial layout of the source video, and dilated dual masking to ensure smooth appearance transfer.
  • We present zero image guidance, enhancing color saturation via controllable negative prompt blending.
  • Comprehensive quantitative and qualitative evaluations demonstrate that our method outperforms existing state-of-the-art approaches in both quality and efficiency.
@@ 2 Methods
We introduce VINO, a training-free video editing approach that performs object replacement while maintaining overall scene coherence using text and image prompts, as illustrated in Fig. 4. Our method is grounded in the careful integration of well-structured noise with appropriately chosen components. Specifically, we consider two key ingredients: the foreground latent, obtained via $\rho$-start sampling, and the background latent, acquired through DDIM inversion of the source video. These latents are blended using a tailored mask derived from both the source and target object masks, employing dilated dual masking to ensure smooth transitions and spatial consistency. In addition, we apply zero image guidance, a negative prompt-based strategy, to enhance color saturation and overall visual fidelity in the final output.
@@ 2.1 Preliminaries
Latent Diffusion Models (LDMs).
LDMs [stablediffusion] perform the diffusion process in a compressed latent space obtained from an autoencoder $\mathcal{E},\mathcal{D}$. Given an input image $\bm{x}$}, the encoder produces a clean latent representation ${\bm{z}}_{0}=\mathcal{E}({\bm{x}})$, where the diffusion process adds and removes noise as:
${\bm{z}}_{t}=\sqrt{\bar{\alpha}_{t}}\,{\bm{z}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,{\mathbf{I}}).$ (1)
To balance realism and prompt adherence, classifier-free guidance [CFG] combines conditional and unconditional diffusion model predictions as follows:
$\hat{\bm{\epsilon}_{\theta}}=(1+w)\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,{\bm{c}})-w\,\bm{\epsilon}_{\theta}({\bm{z}}_{t},t,\bar{{\bm{c}}}),$ (2)
where $w$ is the guidance scale, $\bm{c}$} is the conditioning input, and $\bar{{\bm{c}}}$ denotes the negative conditioning input. This design enables efficient training and sampling while maintaining high-quality visual fidelity consistently.
Figure 5: Controlled noise sampling for structured edits in VINO. Foreground and background latents, obtained via $\rho$-start sampling and zero image guidance, are blended using a final mask formed by the union of dilated source and target masks. This enables clean removal of the source object and seamless insertion of the target.
DDIM Sampling and Inversion.
DDIM [ddim] provide a deterministic and efficient sampling process that reduces the number of steps required for faithful generation. The sampling process updates the latent as:
${\bm{z}}_{t-1}$ $=\sqrt{\alpha_{t-1}}\left(\frac{{\bm{z}}_{t}-\sqrt{1-\alpha_{t}}\,\bm{\epsilon}_{\theta}({\bm{z}}_{t},t)}{\sqrt{\alpha_{t}}}\right)$ (3)
$\quad+\sqrt{1-\alpha_{t-1}}\cdot\bm{\epsilon}_{\theta}({\bm{z}}_{t},t),$
where we denote this update rule as $\text{DDIM-samp}({\bm{z}}_{t},t)$. The inversion process can be obtained by reversing these sampling steps, enabling reconstruction and fine-grained editing in the latent space. For editing, specific regions can be modified by blending DDIM inversion latents from the source and ongoing edit using a spatial mask:
${\bm{z}}^{\prime}_{t}={\bm{z}}^{\text{{fg}}}_{t}\odot{\bm{M}}+{\bm{z}}^{\text{{bg}}}_{t}\odot(1-{\bm{M}}),$ (4)
where $\bm{M}$} specifies the edited region, ${\bm{z}}^{\text{{fg}}}_{t}$ and ${\bm{z}}^{\text{{bg}}}_{t}$ denote background and foreground latents, respectively.
@@ 2.2 Overview
Our approach achieves this goal through a two-stage editing pipeline, as shown in Fig. 4. In the first stage, we construct a rough noise to enable localized object modification. In the second stage, this rough noise is refined into a more well-structured good noise, leading to higher-fidelity editing.
@@ 2.3 How to construct a good noise for video editing?
We first encode the source video $\mathcal{V}_{s}$ into the latent space using a pretrained VAE encoder $\mathcal{E}$. We then initialize the noisy latent using $\rho$-start sampling (Section 2.3.2), a modification of the DDIM sampling schedule, which enables better preservation of the source video’s structure and motion. This noisy latent is passed into a pretrained T2V diffusion model $\epsilon_{\theta}$ where we incorporate both image and text guidance using the IP-Adapter [ipadapter]. The image and text features from CLIP image encoder $\text{CLIP}_{\text{img}}$ and CLIP text encoders $\text{CLIP}_{\text{txt}}$ [clip] are injected into each corresponding cross-attention layers of the T2V model as condition embeddings:
@@
During the denoising process with condition embeddings, we employ a foreground-background separation mechanism to perform object-centric modifications while preserving background fidelity, inspired by latent blending [latentblending, blendeddiffusion, videoswap, genvideo]. Specifically, we split the noisy latent ${\bm{z}}_{t}$ into foreground and background regions using the object segmentation mask $\bm{M}$}. For the background region, we retrieve the DDIM inversion noise corresponding to the current time step to preserve the background of the source video $\mathcal{V}_{s}$. In contrast, the foreground region is denoised by iteratively updating the noisy latent from the previous time step using a denoising network $\epsilon_{\theta}$, guided by the condition embeddings. This enables the foreground object to be progressively edited to resemble the target object. Formally, the foreground latent ${\bm{z}}_{t}^{\text{{fg}}}$ and background latent ${\bm{z}}_{t}^{\text{{bg}}}$ at time step $t$ are computed as:
${\bm{z}}^{\text{{fg}}}_{t\texttt{-}1}$ $=\text{DDIM-samp}({\bm{z}}^{\prime}_{t},{\bm{c}}_{\text{img}},{\bm{c}}_{\text{txt}},\bar{{\bm{c}}}_{\text{img}},\bar{{\bm{c}}}_{\text{txt}},t),$ (6)
${\bm{z}}^{\text{{bg}}}_{t\texttt{-}1}$ $={\bm{z}}_{t\texttt{-}1},\ {\bm{z}}_{t-1}\in\{{\bm{z}}_{t}\}_{t=1}^{\tau_{\nu}}\text{ from }\text{DDIM-inv}({\bm{z}}_{0}).$ (7)
We then blend these two latents using $\bm{M}$} via our dilated dual masking (see Section 2.3.1), as follows:
@@
where $\odot$ denotes element-wise multiplication. We use different mask $M$ at each stage. Specifically, in the first stage, the dilated source mask from the source video, ${\bm{M}}^{\prime}_{s}\in\mathbb{R}^{F\times C_{s}\times H_{s}\times W_{s}}$, is used. Through the denoising process in the first stage, we obtain a roughly edited video $\mathcal{V}_{r}$, which reflects the general structure and appearance of the target object. However, this rough edit may still retain residual features of the source object, which should be removed for a clean visual transformation ultimately.
@@ 2.3.1 Dilated dual masking
Our proposed method coordinates structured noise to define editable regions in each frame and enables selective editing of target objects (Fig. 5). Unlike Blended Latent Diffusion (BLD) [blendeddiffusion], which gradually shrinks masks during diffusion to address resolution degradation, our dilated dual masking uses a fixed dilation to refine boundaries and better spatially align source–target shapes.
In the first stage, only the source object mask ${\bm{M}}_{s}$ is available, which roughly outlines the foreground region to be edited (Fig. 4(b-1)). However, this mask lacks contextual adaptation to the background, often resulting in sharp edges and unnatural transitions at the object boundaries. Thus, we apply morphological dilation to smooth boundaries and facilitate seamless object replacement (Fig. 4(a-1)). Specifically, a single-iteration morphological dilation is applied to a binary matrix $M$ with a kernel of size $k\times k$. For each pixel position $(p,q)$, the output value is determined as the maximum value within a neighborhood defined by ${\mathcal{N}}_{k}$, where ${\mathcal{N}}_{k}=\{(i,j)\mid-r\leq i,j\leq r\}$ and $r=\lfloor\frac{k}{2}\rfloor$. The dilation operation is formally expressed as:
@@
In the second stage, we similarly obtain the dilated rough mask ${\bm{M}}^{\prime}_{r}$ from the rough edits $\mathcal{V}_{s}$ (Fig. 4(b-2)). However, blending foreground and background latents using either the dilated source mask ${\bm{M}}^{\prime}_{s}$ or the dilated rough mask ${\bm{M}}^{\prime}_{r}$ poses the risk that parts of the source video may remain visible or that the generation area for the target object may not be fully covered. To address this, we leverage masks from both the first and second stages (Fig. 4(a-2)). The final mask ${\bm{M}}_{e}$ is computed as the union of the dilated source mask ${\bm{M}}^{\prime}_{s}$ and the dilated rough mask from the first stage ${\bm{M}}^{\prime}_{r}$, as shown in the following equation below:
@@
Finally, we construct the good noise ${\bm{z}}^{\prime}_{t}$ using the final mask ${\bm{M}}_{e}$ to guide the final blending process in Eq. 8. This enables the second stage to generate more realistic, artifact-free results, as ${\bm{M}}_{e}$ precisely defines the edited region. In particular, it ensures clean removal of the original object, accurate reflection of the new object’s shape, and consistent preservation of the background appearance consistently.